OpenStudy (anonymous):

Help!!!! Please 3x^2-6x x-2 ________divided________ 3x-5 3x-5

OpenStudy (anonymous):

\[\frac{ 3x^2-6x }{3x-5} \times \frac{ 3x-5 }{ x-2 }\] then cancel ( 3x-5 ) by ( 3x-5 ) You'll get ( 3x^2-6x ) ( x-2 ) = 3x^3 - 6x^2 - 6x^2 + 12 x = 3x^3 - 12x^2 + 12x = 3x ( x^2 - 4x -4 )

OpenStudy (anonymous):

So you flip the denominator

OpenStudy (anonymous):

yes when you divide by a fraction you flip it upside down & multiply. \[\frac{ a }{ b } \div \frac{ x }{ y } = \frac{ a }{ b } \times \frac{ y }{ x }\]

OpenStudy (anonymous):

Thank You. I am not good at math at all...I had someone helping me before. He was amazing. I just dont get it.

OpenStudy (anonymous):

You're welcome. :) Good luck.